OpenStudy (anonymous):

pefrom the indicated operation. wirte the result in the from atbi: (7+9i)(7-9i)

OpenStudy (anonymous):

corecction a+bi

OpenStudy (watchmath):

Using the identity that \((a+b)(a-b)=a^2-b^2\) we simplify the expression into \(7^2-(9i)^2=49+81=130\)

OpenStudy (anonymous):

\[(7+9i)(7-9i)=49+81=130\]

OpenStudy (anonymous):

a bag of battiries contains 5 batteries which are all the same size and equally selected. each battery is a different brand. if u selected 4 at ramdom. use the wingspaning pricinclple to determine how many points will be in the sample space with or without?

OpenStudy (anonymous):

You beat me to it watchmath :)

OpenStudy (anonymous):

factor by grouping: 6st+21t-10s-35

OpenStudy (anonymous):

\[6st+21t-10s-35=(6st-10s)+(21t-35) =2s(3t-5)+7(3t-5)\] \[=(3t-5)(2s+7)\]

OpenStudy (anonymous):

x=√4x+12 solve √=this is a square root sign

OpenStudy (anonymous):

how do i give medals. id give you a milllion

OpenStudy (anonymous):

Subtract 12 from both sides, and then square both sides.

OpenStudy (anonymous):

Lol. Press on "Good Answer"

OpenStudy (anonymous):

i cant find it!

OpenStudy (anonymous):

\[x=\sqrt{4x}+12 \implies (x-12)^2=4x \implies x^2-24x+144-4x=0\] \[\implies x^2-28x+144=0\] Now, use the quadratic formula to solve the equation.

OpenStudy (anonymous):

found the buttoN! lol

OpenStudy (anonymous):

given f(x)=x^2-4x-6, find f(2)

OpenStudy (anonymous):

\[f(x)=x^2-4x-6 \implies f(2)=2^2-4(2)-6=4-8-6=-10\]

OpenStudy (anonymous):

perform the indicated operation and simplify: 3/t+5/r-6 the division sign means over

OpenStudy (anonymous):

Are you sure about the expression?

OpenStudy (anonymous):

oh sorry its 3/r + 5/r-6

OpenStudy (watchmath):

\(\frac{3(r-6)}{r(r-6)}+\frac{5r}{r(r-6)}=\frac{8r-18}{r(r-6)}\)

OpenStudy (anonymous):

simplify and add or subratct wherever possible: √3-6√12=4√192